If i flipped a fair coin 5 times, what is the probability of getting 3 heads.
Please show work..;)
thank you so much

Answer by stanbon(75887)

If i flipped a fair coin 5 times, what is the probability of getting 3 heads.
Please show work..;)
--------------------------
The 3 heads can happen in 5C3 = 10 ways
The probability of any one of the 3 head/2 tail patterns is (1/2)^3*(1/2)^2
------------
The total probability is 5C3(1/2)^3*(1/2)^2
= 10*(1/2)^5
= 10/32
= 5/16
============
Cheers,
Stan H.